If the equation of a circle is x^2+y^2+3x-my+10=0 and the point (2,-5) lies on the circle, what is the value of m? I think it is -6???
Answer by stanbon(75887)

If the equation of a circle is x^2+y^2+3x-my+10=0 and the point (2,-5) lies on the circle, what is the value of m? I think it is -6???
--------------------------------------
2^2+5^2+3*2+5m+10 = 0
4 + 25 + 6 + 5m + 10 = 0
5m = -45
m = -9
=============
Cheers,
Stan H.